OpenStudy (stormoshawty):

Solve using elimination. 2x + 4y = 22 2x - 2y = -8

OpenStudy (shamil98):

multiply second equation by two, what do u get

OpenStudy (shamil98):

after that you can use elimnation

OpenStudy (stormoshawty):

2x + 4y = 22 4x - 4y = -16?

OpenStudy (anonymous):

now add both equations together

OpenStudy (stormoshawty):

6x + y = 6?

OpenStudy (shamil98):

no. 6x = 6 *

OpenStudy (shamil98):

the y's cancel out

OpenStudy (stormoshawty):

oooh yeah, my bad.

OpenStudy (shamil98):

anyways plug your value of x back into the equation 2x + 4y = 22 and solve for y.

OpenStudy (stormoshawty):

2x + 4y = 22 2(1) + 4y = 22 6y = 22 Y = 3.666666

OpenStudy (stormoshawty):

Right? @shamil98

OpenStudy (anonymous):

No, when you plug in (1): 2(1) + 4 y = 22 2 + 4y = 22 4y = 20 y = 5 Your answers are x = 1 and y = 5

OpenStudy (anonymous):

Remember that you can only add like terms. You cannot add 2 + 4y to get 6y because 2 does not have a y.

OpenStudy (stormoshawty):

yeahhh. I rememer, thank you! :)